OpenStudy (anonymous):

the speed of a freight train is 20 mph slower than the speed of passenger train. the freight train travels 240 miles in the same time that it takes the passenger train to travel 320 miles. find the speed of each train.

OpenStudy (anonymous):

if x= freight train speed y=passenger train speed then: x=y-20 240/x=320/y solve it for x & y

OpenStudy (anonymous):

240y=320(y-20) y=80 x=60